# Let Them Down Easy! Contextual Effects of LLM Guardrails on User Perceptions and Preferences

### v1

@@ A.3 Human Annotation of Query-Motivation Pairs
We recruited 672 annotators in total via Prolific [^7] https://www.prolific.com/ using Potato Pei and Jurgens (2023) to label 280 queries, resulting in 560 query-motivation instances (two motivations per query). Each instance is labeled by 6 annotators and each annotator labels 5 instances. We show the annotation instruction in Figure 11 and show the screenshots of the instruction, consent form, and annotation interface in Figure 9 and 10.
@@ Appendix B LLM Experimental Details and Hyperparameter Setting
Skywork-Reward-Llama-3.1-8B-v0.2 Liu et al. (2024) Skywork Community License
Skywork-Reward-Gemma-2-27B-v0.2 Liu et al. (2024) Skywork Community License
QRM-Llama3.1-8B Dorka (2024) Llama 3 Community License Agreement
QRM-Gemma-2-27B Dorka (2024) Llama 3 Community License Agreement
tulu-v2.5-13b-preference-mix-rm Ivison et al. (2024) Apache license 2.0
Llama-3.3-70B-Instruct Grattafiori et al. (2024) Llama 3.3 Community License Agreement
Mistral-7B-Instruct-v0.3 Jiang et al. (2023) Apache license 2.0
OLMo-2-32B OLMo et al. (2024) Apache license 2.0
Qwen2.5-Instruct (7B, 72B; Team, 2024) qwen license
@@ C.4 User Study Statistical Analysis
   Direct Refusal $-2.36^{***}$
   Explanation-based $-1.45^{***}$
   Partial Compliance $-0.62^{***}$
   Redirection $-1.71^{***}$
@@
Alignment (Misaligned) $-0.38^{***}$ $0.013^{***}$
[…]
@@ C.6 User Study Data Validation
   Direct Refusal $-4.18$ 0.08 $<$.001
   Explanation-based $-3.39$ 0.08 $<$.001
   Redirection $-3.13$ 0.08 $<$.001
   Partial Compliance $-1.34$ 0.08 $<$.001
@@
   Privacy Violations $0.31$ 0.07 $<$.001
   Triggers for Offensive Language $-0.18$ 0.07 .013
Query Status (Malicious) $-0.09$ 0.05 .077 0.0008
Alignment (Misaligned) $-0.01$ 0.05 .905 $<$0.0001
[…]
@@ C.7 Response Strategies Impact
Intercept $5.38^{***}$ $5.96^{***}$ $5.52^{***}$ $5.49^{***}$ $5.35^{***}$ $2.30^{***}$ $1.88^{***}$
Direct Refusal $-3.68^{***}$ $-2.36^{***}$ $-3.45^{***}$ $-2.98^{***}$ $-2.63^{***}$ $2.85^{***}$ $1.58^{***}$
Explanation-based Refusal $-3.08^{***}$ $-1.45^{***}$ $-2.68^{***}$ $-2.19^{***}$ $-1.10^{***}$ $2.31^{***}$ $1.20^{***}$
Partial Compliance $-1.42^{***}$ $-0.62^{***}$ $-1.41^{***}$ $-1.21^{***}$ $-0.94^{***}$ $1.11^{***}$ $0.26^{**}$
Redirection $-2.79^{***}$ $-1.71^{***}$ $-2.74^{***}$ $-2.25^{***}$ $-1.77^{***}$ $2.27^{***}$ $1.10^{***}$
[…]
@@
Main Effects Direct Refusal $-3.25^{***}$ $-2.41^{***}$ $-3.14^{***}$ $-2.73^{***}$ $-2.49^{***}$ $2.58^{***}$ $1.48^{***}$
Explanation-based $-2.70^{***}$ $-1.28^{***}$ $-2.36^{***}$ $-1.91^{***}$ $-0.99^{***}$ $1.97^{***}$ $1.06^{***}$
Partial Compliance $-0.76^{***}$ $-0.02$ $-0.76^{***}$ $-0.69^{***}$ $-0.44^{***}$ $0.61^{***}$ $-0.03$
Redirection $-2.26^{***}$ $-1.38^{***}$ $-2.21^{***}$ $-1.82^{***}$ $-1.41^{***}$ $1.87^{***}$ $0.91^{***}$
Malicious Status $0.17^{**}$ $0.33^{***}$ $0.47^{***}$ $0.41^{***}$ $0.35^{***}$ $-0.34^{***}$ $-0.15^{**}$
Misaligned $-0.80^{***}$ $-0.64^{***}$ $-0.80^{***}$ $-0.74^{***}$ $-0.51^{***}$ $0.69^{***}$ $0.34^{***}$
Interactions Direct × Malicious $-0.85^{***}$ $0.11$ $-0.61^{***}$ $-0.50^{***}$ $-0.28^{.}$ $0.54^{***}$ $0.19$
Explanation × Malicious $-0.76^{***}$ $-0.33^{*}$ $-0.65^{***}$ $-0.56^{***}$ $-0.23$ $0.69^{***}$ $0.29^{*}$
Partial × Malicious $-1.32^{***}$ $-1.18^{***}$ $-1.29^{***}$ $-1.06^{***}$ $-1.00^{***}$ $0.99^{***}$ $0.59^{***}$
Redirection × Malicious $-1.06^{***}$ $-0.67^{***}$ $-1.06^{***}$ $-0.88^{***}$ $-0.72^{***}$ $0.79^{***}$ $0.39^{**}$
[…]

### v2

@@ A.3 Human Annotation of Query-Motivation Pairs
We recruited 672 annotators in total via Prolific [^7] https://www.prolific.com/ using Potato pei2023annotator to label 280 queries, resulting in 560 query-motivation instances (two motivations per query). Each instance is labeled by 6 annotators and each annotator labels 5 instances. We show the annotation instruction in Figure 11 and show the screenshots of the instruction, consent form, and annotation interface in Figure 9 and 10.
@@ Appendix B LLM Experimental Details and Hyperparameter Setting
Skywork-Reward-Llama-3.1-8B-v0.2 liu2024skywork Skywork Community License
Skywork-Reward-Gemma-2-27B-v0.2 liu2024skywork Skywork Community License
QRM-Llama3.1-8B dorka2024quantile Llama 3 Community License Agreement
QRM-Gemma-2-27B dorka2024quantile Llama 3 Community License Agreement
tulu-v2.5-13b-preference-mix-rm ivison2024unpacking Apache license 2.0
Llama-3.3-70B-Instruct grattafiori2024llama Llama 3.3 Community License Agreement
Mistral-7B-Instruct-v0.3 jiang2023mistral7b Apache license 2.0
OLMo-2-32B olmo20242 Apache license 2.0
Qwen2.5-Instruct (7B, 72B; qwen2.5) qwen license
@@ C.4 User Study Statistical Analysis
   Direct Refusal $2.36^{***}$
   Explanation-based $1.45^{***}$
   Partial Compliance $0.62^{***}$
   Redirection $1.71^{***}$
@@
Alignment (Misaligned) $38^{***} 0.0$$013^{***}$
Model: $R^{2}=0.230$, $F(9,3830)=127.2$, $p<.001$
Table 10: OLS regression predicting ethical appropriateness from response strategy, query status, guardrail alignment, and query category. Response strategy dominates ethical judgments ($\eta^{2}=0.219$), showing that what models do strongly shapes what users believe models should do, far exceeding the influence of user intent ($\eta^{2}=0.0004$). * $p<.05$, ** $p<.01$, *** $p<.001$
@@ C.6 User Study Data Validation
   Direct Refusal 4.18-4.18 0.08 <<.001
   Explanation-based 3.39-3.39 0.08 <<.001
   Redirection 3.13-3.13 0.08 <<.001
   Partial Compliance 1.34-1.34 0.08 <<.001
@@
   Privacy Violations $0.31$ 0.07 <<.001
   Triggers for Offensive Language 0.18-0.18 0.07 .013
Query Status (Malicious) 0.09-0.09 0.05 .077 0.0008
Alignment (Misaligned) 0.01-0.01 0.05 .905 <<0.0001
Model: $R^{2}=0.519$, $F(9,3830)=459.4$, $p<.001$
Table 12: OLS regression predicting perceived compliance from response strategy, query status, guardrail alignment, and query category. Response strategy explains the majority of variance ($\eta^{2}=0.516$), while other predictors have minimal impact. *** $p<.001$, * $p<.05$
@@ C.7 Response Strategies Impact
Intercept $5.38^{***}$ $96^{***} 5.5$$52^{***}$ $5.49^{***}$ $5.35^{***}$ $2.30^{***}$ $1.88^{***}$
Direct Refusal $3.68^{***}$ $2.36^{***}$ $3.45^{***}$ $2.98^{***}$ $2.63^{***}$ $2.85^{***}$ $1.58^{***}$
Explanation-based Refusal $3.08^{***}$ $1.45^{***}$ $2.68^{***}$ $2.19^{***}$ $1.10^{***}$ $2.31^{***}$ $1.20^{***}$
Partial Compliance $1.42^{***}$ $0.62^{***}$ $1.41^{***}$ $1.21^{***}$ $0.94^{***}$ $1.11^{***}$ $0.26^{**}$
Redirection $2.79^{***}$ $1.71^{***}$ $74^{***} 2.2$$25^{***}$ $1.77^{***}$ $2.27^{***}$ $1.10^{***}$
Table 13: OLS regression coefficients showing the effect of each refusal strategy on user perceptions relative to Full Compliance. All coefficients are unstandardized. Significance levels: $^{.}p<0.1$$^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$. Direct refusals consistently produce the most negative reactions across dimensions, while partial compliance shows the least negative impact.
@@
Main Effects Direct Refusal $3.25^{***}$ $2.41^{***}$ $3.14^{***}$ $2.73^{***}$ $2.49^{***}$ $2.58^{***}$ $1.48^{***}$
Explanation-based $2.70^{***}$ $1.28^{***}$ $2.36^{***}$ $1.91^{***}$ $0.99^{***}$ $1.97^{***}$ $1.06^{***}$
Partial Compliance $76^{***} 0.0$2-0.02 $0.76^{***}$ $0.69^{***}$ $0.44^{***}$ $61^{***} 0.0$3-0.03
Redirection $2.26^{***}$ $1.38^{***}$ $2.21^{***}$ $1.82^{***}$ $1.41^{***}$ $1.87^{***}$ $0.91^{***}$
Malicious Status $0.17^{**}$ $0.33^{***}$ $0.47^{***}$ $0.41^{***}$ $0.35^{***}$ $0.34^{***}$ $0.15^{**}$
Misaligned $0.80^{***}$ $0.64^{***}$ $0.80^{***}$ $0.74^{***}$ $0.51^{***}$ $0.69^{***}$ $0.34^{***}$
Interactions Direct × Malicious $0.85^{***}$ $0.11$ $0.61^{***}$ $0.50^{***}$ $0.28^{.}$ $0.54^{***}$ $0.19$
Explanation × Malicious $0.76^{***}$ $0.33^{*}$ $^{***} 0.56-0.56^{***}$ 0.23-0.23 $0.69^{***}$ $0.29^{*}$
Partial × Malicious $32^{***} 1.1$$18^{***}$ $1.29^{***}$ $6^{***} 1.00-1$$0^{***}$ $0.99^{***}$ $0.59^{***}$
Redirection × Malicious $1.06^{***}$ $0.67^{***}$ $1.06^{***}$ $0.88^{***}$ $0.72^{***}$ $0.79^{***}$ $0.39^{**}$
Table 14: OLS regression results for predicting user perceptions from response strategy, query intent, and guardrail alignment. The model includes response strategy × status interactions and guardrail alignment as an independent main effect. Significance levels: $^{.}p<0.1$$^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$.